OpenStudy (anonymous):

1. Use the Substitution Method to solve the following system of equations. x - 4y = 1 2x - 9y = 3 (-1, -3) (-3, 1) (3, 1) (-3, -1) 2. What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 1 3x - 4y = 10 2 -1 1 -2

OpenStudy (anonymous):

x=4y+1 2x-9y=3

OpenStudy (anonymous):

2(4y+1)-9y=3 solve for y

OpenStudy (anonymous):

is it (3,1)

OpenStudy (anonymous):

not quite. look at the equation i put up and solve it again

OpenStudy (anonymous):

for 1) is (-3,-1) and for 2 ) is 2 that is the right answer

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

can you help me with some more please

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

1. What is the value of the y variable in the solution to the following system of equations? 5x + 3y = 7 3x - 5y = -23 -1 1 -4 4 2. For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12 -2y - 4 2y - 4 2y + 4 -2y + 4

OpenStudy (anonymous):

The graph below is represented by which system of inequalities? -4x + 2y less than or equal to 8 y < -x + 4 -4x + 2y < 8 y less than or equal to -x + 4 -4x + 2y greater than or equal to 8 y > -x + 4 -4x + 2y greater than or equal to 8 y < -x + 4

OpenStudy (anonymous):

for ur first one the y is 4 and for 2 the answer i got is x=4 and y=0